# Skin Effect Induced Anomalous Dynamics from Charge-Fluctuating Initial States

### v1

@@ Abstract
Introduction.—The study of non-equilibrium dynamics, especially in open systems, has recently emerged as a major research frontier, revealing intriguing phenomena beyond the scope of conventional equilibrium descriptions Diehl et al. (2008, 2011); Barontini et al. (2013); Dast et al. (2014); Huang et al. (2018); Li et al. (2019); Jian et al. (2020); Lapp et al. (2019); Fuji and Ashida (2020); Choi et al. (2020); Huang and Liu (2020); Takasu et al. (2020); Gullans and Huse (2020); Bao et al. (2020); Lu and Grover (2021); Sang and Hsieh (2021); Turkeshi et al. (2021); Bätge et al. (2021); Wald et al. (2021); Levy et al. (2021); Li et al. (2021); Turkeshi and Schiró (2023); Kawabata et al. (2023); Moreira et al. (2024); Paz et al. (2024); Lessa et al. (2025); Despres (2025); Gao and Muttalib (2025); Glatthard (2025); Capecelatro et al. (2025). Non-Hermitian Hamiltonians offer an effective theoretical framework for describing open quantum systems Ashida et al. (2020); Prosen (2008); Lieu et al. (2020); Prosen (2010); Zhou et al. (2020); Zhou and Yu (2022); Ren et al. (2022a). Such non-Hermiticity naturally emerges when a system couples to an environment, introducing dissipation, gain, or non-reciprocal interactions. Moreover, non-Hermitian Hamiltonians provide a conceptually valuable description for quasiparticles possessing finite lifetimes due to mechanisms like electron-phonon coupling or disorder scatteringShen and Fu (2018); Nagai et al. (2020); Huang et al. (2025). The investigation of non-Hermitian systems has unveiled a wealth of exotic phenomena lacking counterparts in Hermitian systems Rudner and Levitov (2009); Hu and Hughes (2011); Esaki et al. (2011); Heiss (2012); Leykam et al. (2017); Couvreur et al. (2017); Nakagawa et al. (2018); Gong et al. (2018); McDonald et al. (2018); Xiong (2018); El-Ganainy et al. (2018); Martinez Alvarez et al. (2018); Kawabata et al. (2019a); Liu et al. (2019); Lee et al. (2019a); Yamamoto et al. (2019); Pan et al. (2020); Zhai et al. (2020); Iskin (2021); Lv et al. (2022); Kawabata et al. (2021); Zhai et al. (2022); Liang et al. (2022); Lee and Thomale (2019); Lee et al. (2019b); Chang et al. (2020); Li et al. (2020); Bergholtz et al. (2021); Xi et al. (2021); Ding et al. (2022); Kozii and Fu (2024); Guo et al. (2023); Shen et al. (2023); Okuma and Sato (2023); Mao et al. (2024); Zhong et al. (2024); Guo et al. (2020); Yang et al. (2020); Sun et al. (2020); Hu and Zhao (2021); Li and Hu (2023); Zheng et al. (2024); Ding et al. (2023); Chen et al. (2024); Gao et al. (2022); Chen et al. (2022); Zhang et al. (2023a); Wang et al. (2022); Zhang et al. (2022a); Yang et al. (2024); Hu et al. (2023); Jing et al. (2024); Yu et al. (2024); Murshed and Roy (2024); Juričić and Roy (2024); Tang et al. (2024); Liu and Batchelor (2025) such as the celebrated non-Hermitian skin effect (NHSE) with macroscopic accumulation of eigenstates localized at the system’s edges Yao and Wang (2018); Yokomizo and Murakami (2019); Yao et al. (2018); Kawabata et al. (2019b); Yokomizo and Murakami (2019); Okuma et al. (2020); Zhang et al. (2020); Song et al. (2019); Helbig et al. (2020); Kawabata et al. (2020); Zou et al. (2021); Zhou et al. (2022); Longhi (2021); Kim and Park (2021); Yokomizo and Murakami (2021); Mao et al. (2021); Xiao et al. (2020); Zhang et al. (2022b); Guo et al. (2022); Feng et al. (2025); Sun et al. (2021); Li et al. (2023a); Yoshida et al. (2024); Huang et al. (2024); Wang et al. (2024a); Zhang et al. (2024); Hu (2025); Sun et al. (2025).
@@
While significant progress has been made in understanding non-Hermitian dynamics, much of the focus has been on systems with particle-number conservation. However, many physical processes particularly those involving Cooper pairing and superconductivity, inherently lead to states that lack a definite particle number, breaking U(1) symmetry. The non-equilibrium dynamics of such U(1)-breaking states under non-Hermitian evolution, and the interplay between particle number fluctuations and unique non-Hermitian features like the NHSE, remain largely unexplored. This regime presents fundamental challenges and opportunities to discover novel phenomena arising from the confluence of non-unitary dynamics, non-reciprocity, and symmetry breaking.
@@
Probing the non-equilibrium evolution of quantum many-body states requires analyzing multiple facets of their properties. Fundamental quantities like particle density provide essential information about the spatial distribution and transport of particles, crucial for understanding phenomena like NHSE. Beyond density, quantum entanglement encapsulates the unique non-classical correlations fundamental to quantum systems Eisert et al. (2010); Calabrese and Cardy (2004); Alba and Calabrese (2017); Ashida and Ueda (2018); Skinner et al. (2019); Herviou et al. (2019); Chen et al. (2020); Bácsi and Dóra (2021); Orito and Imura (2022); Dóra et al. (2022); Turkeshi and Schiró (2023); Kawabata et al. (2023); Orito and Imura (2023); Li et al. (2023b, 2024). Monitoring its evolution, often quantified as entanglement entropy (EE), provides a valuable dynamic window into a system’s properties Cao et al. (2019); Fuji and Ashida (2020); Alberton et al. (2021); Turkeshi et al. (2021); Kells et al. (2023); Wang et al. (2024b). Furthermore, to specifically track the restoration of broken U(1) symmetry, entanglement asymmetry (EA) has been introduced as a sensitive probe derived from entanglement, capable of detecting subtle changes in the symmetry properties of the state  Ares et al. (2023a). Analyzing EA dynamics in Hermitian systems has revealed intriguing symmetry restoration phenomena, including the quantum Mpemba effect (QME) Nava and Fabrizio (2019); Carollo et al. (2021); Zhang et al. (2025); Moroder et al. (2024); Bettmann and Goold (2025); Chatterjee et al. (2023, 2024); Aharony Shapira et al. (2024); Manikandan (2021); Kochsiek et al. (2022); Ivander et al. (2023); Zhou et al. (2023); Wang and Wang (2024); Liu et al. (2024a); Longhi (2024); Nava and Egger (2024); Boubakour et al. (2025); Foligno et al. (2025); Zatsarynna et al. (2025); Longhi (2025); Dong et al. (2025); Graf et al. (2025); Van Vu and Hayakawa (2025); Teza et al. (2025); Joshi et al. (2024); Chang et al. (2024); Caceffo et al. (2024); Strachan et al. (2024), where states initially exhibiting stronger symmetry breaking can paradoxically restore symmetry faster Ares et al. (2023a, b); Murciano et al. (2024); Chalas et al. (2024); Yamashika et al. (2024a); Rylands et al. (2024a); Yamashika et al. (2024b); Rylands et al. (2024b); Liu et al. (2024b, c); Yu et al. (2025); Giulio et al. (2025); Klobas et al. (2025); Ares et al. (2025); Ferro et al. (2024); Klobas (2024). While these quantities have been extensively studied in various contexts, their interplay and behavior under non-Hermitian evolution, particularly for initial states lacking particle-number conservation, remain elusive. This motivates fundamental questions: How are particle density profiles, entanglement growth, and U(1) symmetry restoration affected by non-Hermitian dynamics and the NHSE in states without definite particle number?
@@
To address these essential questions, we investigate the non-equilibrium dynamics of a superconducting pairing state under the time evolution of a typical non-Hermitian Hamiltonian, termed the Hatano-Nelson (HN) model Hatano and Nelson (1997), which involves non-reciprocal hopping of fermions exhibiting the skin effect. For the time evolution of the pairing state under HN model, the total particle number is not conserved. Here, we demonstrate a universal late time convergence to half-filling, arising from the skin effect, regardless of the initial particle number. Furthermore, we uncover a range of novel phenomena in the short-time dynamics arising from the interplay of U(1) symmetry breaking and non-Hermiticity. (1) We observe a chiral wavefront in the particle density dynamics, which is also manifested in the behavior of particle inflow due to non-Hermiticity for certain initial states. (2) The rate of entanglement growth is enhanced at early times while greatly suppressed at late times by the non-Hermiticity. (3) In the process of U(1) symmetry restoration, QME persists with the introduction of non-Hermiticity. Crucially, these universal results are robust across a wide range of parameters within the model, holding significant promise for experimental verification.
@@
where $\hat{c}_{j}^{\dagger}$ ($\hat{c}_{j}$) represent the creation (annihilation) operators, $J$ is the amplitude of nearest-neighbor (NN) hopping, $\mu$ is chemical potential, and $\Delta$ denotes the pairing amplitude. The initial state is thus a superconducting BCS pairing state breaking charge U(1) symmetry. Through Jordan-Wigner transformation (see the section I of the supplemrental material (SM) Ref for the details), the fermionic Hamiltonian in Eq. (1) is straightforwardly transformed to a spin-1/2 chain as follows:
@@
$\Delta$ breaks U(1) spin rotational symmetry around $z$-axis. We implement anti-periodic boundary condition (APBC) of Eq. (1), corresponding the periodic boundary condition (PBC) of spin-1/2 chain in Eq. (2) under Jordan-Wigner transformation for even number of lattice sites. The ground state of Eq. (1) is conveniently expressed in the spin representation as a superposition of tilted ferromagnetic states Ares et al. (2023a):
@@
[…]
where $\theta$ represents the tilted angle characterizing the strength of U(1) symmetry breaking. The relations between $\theta$, $\Delta$ and $\mu$ satisfy Ares et al. (2023a, b)
$\text{cos}^{2}\theta=\frac{1-\Delta}{1+\Delta},\Delta^{2}+\mu^{2}=1.$ (4)
To investigate the non-Hermitian dynamics, we employ the HN model Hatano and Nelson (1997) whose Hamiltonian is:
@@
For numerical calculations, we use the quantum software TensorCircuit-NG *[][.https://github.com/tensorcircuit/tensorcircuit-ng.]Zhang2022_z. The energy unit is set as $J=1$, system size $L=64$ and OBC are employed. Without loss of generality, we focus on $\gamma\geq 0$, which results in the skin effect where particles accumulate at the left boundary. It is important to note that the combination of non-reciprocity and the large system size can introduce significant numerical errors in the direct calculation of $e^{-i\hat{H}_{\text{evo}}t}$ Feng et al. (2025). To maintain numerical accuracy, we utilize a small time-step decomposition of the evolution and perform calculations with a high precision of 250 digits (See the details in section II of the SM Ref ).
@@
Figure 2: Entanglement entropy under open boundary conditions. (a), (b) and (c) correspond to $\gamma=0.0$, 0.2 and 0.6, respectively, where the legends represent different tilted angles. The traced subsystem $B$ with length $l=6$ is located at the left end of the system. The total system size is $L=64$.
@@
where $\hat{n}_{j}=\hat{c}^{\dagger}_{j}\hat{c}_{j}$ is the on-site particle density. The results of $\overline{N}(t)$ for the tilted ferromagnetic state with different $\theta$ and fixed $\gamma=0.8$, shown in Fig. 1(a), decrease with time. Most notably, we uncover a universal long-time behavior: regardless of their initial particle number, all tilted ferromagnetic states ($\theta>0$) converge to a total particle number close to the half filling ($L/2$) under long-time evolution. We provide further justification for this observation, including quantitative analysis and a rigorous proof in section III of the SM Ref . Specifically, we rigorously demonstrate that in the limit $\gamma\rightarrow 1$ for this HN model, the total particle number asymptotically approaches $L/2$ as $t\to\infty$. In contrast, the ferromagnetic state ($\theta=0$), as an eigenstate of the on-site particle density operator $\hat{n}_{j}$, remains fully occupied during the evolution. In terms of systems with PBC, the total particle number for a general initial state also tends towards $L/2$ in the steady state.
@@
Density dynamics.— To further explore the dynamical characteristics reflected in the particle number, we now examine the time evolution of the on-site particle density $\langle\hat{n}_{j}\rangle$, exemplified in Fig. 1(b) for $\theta=\pi/6$ Ref . While the initial state exhibits a uniform particle density, distinct behaviors emerge at different timescales. At long times, there is a clear spatial separation of particles due to the skin effect. The particle density becomes concentrated on the left side of the system, leaving the opposite side sparsely populated. Intriguingly, short-time dynamics reveal a contrasting exotic behavior. Instead of accumulation, we observe an initial reduction of particle density near the left edge. This sharp depletion is linked to a chiral wavefront propagating with twice the maximum group velocity of quasiparticles. Specifically, particle density begins to decrease as the wavefront reaches a given site. Here, the group velocity is defined as $v_{\text{g}}=\frac{\mathrm{d}}{\mathrm{d}k}\text{Re}[E(k)]$, where $E(k)=-2J\text{cos}k-2i\gamma\text{sin}k$ is the energy spectrum of the HN model under PBC Modak et al. (2020); Orito and Imura (2023); Cevolani et al. (2018).
@@
With increasing time, two characteristic timescales, $\tau_{1}$ and $\tau_{2}$, become evident. The first $\tau_{1}$ corresponds to the arrival time of the chiral wavefront at the right boundary, at which point the total particle number approaches $L/2$ for large $\gamma$. Around this time, the particle density begins to change from both boundaries inwards: increasing from the left edge and decreasing from the right edge towards the center. Upon reaching the second timescale, $\tau_{2}$, the density distribution stabilizes into a steady state, with only minor oscillations remaining in the central region. This steady state distribution clearly demonstrates the skin effect, with the left half of the system almost fully occupied and the right half nearly empty. In stark contrast, an initial tilted antiferromagnetic (AFM) state lacks this chiral wavefront; its dynamics immediately resemble the post-$\tau_{2}$ behavior of the tilted ferromagnetic state, likely because the AFM state starts at half-filling. Detailed results are presented in section V of the SM Ref .
@@
Similar to the particle density, characteristic timescales $\tau_{1}$ and $\tau_{2}$ are also evident in the dynamics of both current density and particle inflow. In the short-time regime ($\tau<\tau_{1}$), a chiral wavefront emerges in both the current density and particle inflow profiles. The characteristic time $\tau_{1}$ corresponds to the moment when this chiral wavefront reaches the right boundary of the system. Notably, the wavefront in the particle inflow in Fig. 1(d) represents a significant particle sink, indicating a substantial loss of particle number at given space-time location. The wavefront of current in Fig. 1(c) appears as a sharp spatial gradient of current density. Although the negative current gradient implies particle accumulation according to the continuity equation Eq. (8), the observed net effect at the wavefront is particle depletion. Thus, the particle number reduction associated with the wavefront in Fig. 1(b) arises from the effect of particle inflow acting as a propagating sink, which dominates over the particle-accumulating effect of the current induced by non-reciprocity. In the intermediate stage ($\tau_{1}<\tau<\tau_{2}$), the rapid increase in particle number at the left edge and decrease at the right edge are also driven by the particle inflow, as visualized in Fig. 1(d). For late times ($\tau>\tau_{2}$), both the current and particle inflow diminish, exhibiting only small, time-decaying oscillations. Results for other values of $\gamma$ show qualitatively consistent behaviors, as included in section V of the SM Ref . Moreover, our analysis of the density dynamics under APBC, detailed in the the section VI of the SM Ref , yields qualitatively distinct conclusions, further demonstrating that the anomalous phenomena observed under OBC are attributed to the skin effect.
@@
Here, the system is bipartitioned into subsystems $A$ and $B$, and $\rho_{A}=\text{Tr}_{B}[\rho]$ is the reduced density matrix of subsystem $A$. For the Hermitian case (Fig. 2(a)), the EE for different initial tilted states initially increases and subsequently saturates to a steady-state value. In contrast, the non-Hermitian dynamics, shown in Fig. 2(b) and (c), exhibit more complex phenomena. Similar to the particle density dynamics, the entanglement evolution displays distinct behaviors in different temporal regimes. In the stage $\tau<\tau_{1}$, EE increases with evolution time and subsequently saturates, qualitatively consistent with the conventional behavior of entanglement dynamics in Hermitian systems. This trend is qualitatively captured by the quasiparticle picture, as demonstrated in section IV of the SM Ref . An intriguing observation, illustrated in Fig. 2(b)-(c), is that non-Hermiticity can enhance both the growth rate and the maximum value of EE during time evolution for specific initial states and small $\gamma$.
@@
Given the skin effect, we anticipate that the EE depends on the bipartition subsystem’s location. To investigate this, we calculate the EE for different subsystem positions, with detailed results presented in section V of the SM Ref . When the small subsystem is located centrally, the qualitative behavior of EE dynamics remains similar to the cases where the subsystem is at the left or right boundary. However, the reduction in EE after $t>\tau_{1}$ is significantly less pronounced compared to the results in Fig. 2(b) and (c). This milder reduction can be naturally understood through the particle number distribution induced by the skin effect, as the particle density fluctuation is stronger at the middle region.
@@
Recent research has revealed a novel and universal QME in dynamic symmetry restoration, where systems with greater initial symmetry breaking recover symmetry faster Ares et al. (2023a); Murciano et al. (2024); Caceffo et al. (2024); Giulio et al. (2025); Ares et al. (2023b); Chalas et al. (2024); Yamashika et al. (2024a); Rylands et al. (2024a); Yamashika et al. (2024b); Rylands et al. (2024b); Liu et al. (2024b, c); Yu et al. (2025); Chang et al. (2024). A key question is whether symmetry restoration and the QME can emerge in non-Hermitian systems. Fig. 3 presents the second-order Rényi EA for varying non-Hermiticity strengths ($\gamma$). Consistent with prior work Ares et al. (2023a), Fig. 3(a) shows a clear curve crossing for different tilted angles ($\theta$) in the Hermitian case, confirming the QME. Notably, a pronounced QME also exists in the non-Hermitian scenario, as evidenced by curve crossings in Fig. 3(b) and (c). Intriguingly, this QME appears enhanced compared to the Hermitian case, with curve crossings occurring earlier in time. This enhancement is accompanied by non-monotonic behavior in the early-time EA evolution for some initial states, where EA initially increases before decreasing towards symmetry restoration. At late times, the EA results show a restoration of U(1) symmetry, mirroring the behavior observed in Hermitian systems.

### v2

@@ Abstract
Introduction.—The study of non-equilibrium dynamics, especially in open systems, has recently emerged as a major research frontier, revealing intriguing phenomena beyond the scope of conventional equilibrium descriptions [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. Non-Hermitian Hamiltonians offer an effective theoretical framework for describing open quantum systems [31, 32, 33, 34, 35, 36, 37]. Such non-Hermiticity naturally emerges when a system couples to an environment, introducing dissipation, gain, or non-reciprocal interactions. Moreover, non-Hermitian Hamiltonians provide a conceptually valuable description for quasiparticles possessing finite lifetimes due to mechanisms like electron-phonon coupling or disorder scattering[38, 39, 40]. The investigation of non-Hermitian systems has unveiled a wealth of exotic phenomena lacking counterparts in Hermitian systems [41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102] such as the celebrated non-Hermitian skin effect (NHSE) with macroscopic accumulation of eigenstates localized at the system’s edges [103, 104, 105, 106, 107, 104, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128, 129, 130, 131, 132].
@@
While significant progress has been made in understanding non-Hermitian dynamics, much of the focus has been on systems with particle-number conservation [110, 133, 111, 134, 21, 22, 65, 41, 64, 112, 135, 23, 136]. However, many physical processes particularly those involving Cooper pairing and superconductivity, inherently lead to states that lack a definite particle number, breaking U(1) symmetry. The non-equilibrium dynamics of such U(1)-breaking states under non-Hermitian evolution, and the interplay between particle number fluctuations and unique non-Hermitian features like the NHSE, remain largely unexplored. This regime presents fundamental challenges and opportunities to discover novel phenomena arising from the confluence of non-unitary dynamics, non-reciprocity, and symmetry breaking.
@@
Probing the non-equilibrium evolution of quantum many-body states requires analyzing multiple facets of their properties. Fundamental quantities like particle density provide essential information about the spatial distribution and transport of particles, crucial for understanding phenomena like NHSE. Beyond density, quantum entanglement encapsulates the unique non-classical correlations fundamental to quantum systems [137, 138, 139, 136, 140, 141, 142, 143, 144, 145, 22, 23, 146, 147, 148, 149]. Monitoring its evolution, often quantified as entanglement entropy (EE), provides a valuable dynamic window into a system’s properties [150, 9, 151, 152, 17, 153, 154, 155]. Furthermore, to specifically track the restoration of broken U(1) symmetry, entanglement asymmetry (EA) has been introduced as a sensitive probe derived from entanglement, capable of detecting subtle changes in the symmetry properties of the state  [156]. Analyzing EA dynamics in Hermitian systems has revealed intriguing symmetry restoration phenomena, including the quantum Mpemba effect (QME) [157, 158, 159, 160, 161, 162, 163, 164, 165, 166, 167, 168, 169, 170, 171, 172, 173, 174, 175, 176, 177, 178, 179, 180, 181, 182, 183, 184], where states initially exhibiting stronger symmetry breaking can paradoxically restore symmetry faster [156, 185, 186, 187, 188, 189, 190, 191, 192, 193, 194, 195, 196, 197, 198, 199]. While these quantities have been extensively studied in various contexts, their interplay and behavior under non-Hermitian evolution, particularly for initial states lacking particle-number conservation, remain elusive. This motivates fundamental questions: How are particle density profiles, entanglement growth, and U(1) symmetry restoration affected by non-Hermitian dynamics and the NHSE in states without definite particle number?
@@
To address these essential questions, we investigate the non-equilibrium dynamics of a superconducting pairing state under the time evolution of a typical non-Hermitian Hamiltonian, termed the Hatano-Nelson (HN) model [200], which involves non-reciprocal hopping of fermions exhibiting the skin effect. This choice of a U(1) symmetry-breaking initial state marks a significant departure from the aforementioned studies that focused on number-conserving initial states. Our approach therefore provides a direct window into the unique phenomena emerging from the interplay between initial particle number fluctuations and non-Hermitian dynamics. For the time evolution of the pairing state under HN model, the total particle number is not conserved. Here, we demonstrate a universal late time convergence to half-filling, arising from the skin effect, regardless of the initial particle number. Furthermore, we uncover a range of novel phenomena in the short-time dynamics arising from the interplay of U(1) symmetry breaking and non-Hermiticity. (1) We observe a chiral wavefront in the particle density dynamics, which is also manifested in the behavior of particle inflow due to non-Hermiticity for certain initial states. (2) The rate of entanglement growth is enhanced at early times while greatly suppressed at late times by the non-Hermiticity. (3) In the process of U(1) symmetry restoration, QME persists with the introduction of non-Hermiticity. Crucially, these universal results are robust across a wide range of parameters within the model, holding significant promise for experimental verification.
@@
where $\hat{c}_{j}^{\dagger}$ ($\hat{c}_{j}$) represent the creation (annihilation) operators, $J$ is the amplitude of nearest-neighbor (NN) hopping, $\mu$ is chemical potential, and $\Delta$ denotes the pairing amplitude. The initial state is thus a superconducting BCS pairing state breaking charge U(1) symmetry. Through Jordan-Wigner transformation (see the section I of the supplemrental material (SM) [201] for the details), the fermionic Hamiltonian in Eq. (1) is straightforwardly transformed to a spin-1/2 chain as follows:
@@
$\Delta$ breaks U(1) spin rotational symmetry around $z$-axis. We implement anti-periodic boundary condition (APBC) of Eq. (1), corresponding the periodic boundary condition (PBC) of spin-1/2 chain in Eq. (2) under Jordan-Wigner transformation for even number of lattice sites. The ground state of Eq. (1) is conveniently expressed in the spin representation as a superposition of tilted ferromagnetic states [156]:
@@
$|F,\theta\rangle=$ $\underset{j}{\otimes}(\cos\frac{\theta}{2}|\uparrow_{j}\rangle+\sin\frac{\theta}{2}|\downarrow_{j}\rangle),$
where $\theta$ represents the tilted angle characterizing the strength of U(1) symmetry breaking. The relations between $\theta$, $\Delta$ and $\mu$ satisfy [156, 185, 202]
$\text{cos}^{2}\theta=\frac{1-\Delta}{1+\Delta},\Delta^{2}+(\mu/2)^{2}=1.$ (4)
This relation is not an arbitrary constraint but is the necessary condition for our initial state to be a simple, product-form ground state of a local XY Hamiltonian [202]. This allows us to parameterize the initial degree of U(1) symmetry breaking with the single angle $\theta$, a method also employed in Refs. [156, 185] to study Hermitian quench dynamics from such states. To investigate the non-Hermitian dynamics, we employ the HN model [200] whose Hamiltonian is:
@@
For numerical calculations, we use the quantum software TensorCircuit-NG [*[][.https://github.com/tensorcircuit/tensorcircuit-ng.]Zhang2022_z]. The energy unit is set as $J=1$, system size $L=64$ and OBC are employed. Without loss of generality, we focus on $\gamma\geq 0$, which results in the skin effect where particles accumulate at the left boundary. It is important to note that the combination of non-reciprocity and the large system size can introduce significant numerical errors in the direct calculation of $e^{-i\hat{H}_{\text{evo}}t}$ [122]. To maintain numerical accuracy, we utilize a small time-step decomposition of the evolution and perform calculations with a high precision of 250 digits (See the details in section II of the SM [201]).
@@
Figure 2: Entanglement entropy under open boundary conditions. (a), (b) and (c) correspond to $\gamma=0.0$, 0.2 and 0.6, respectively, where the legends represent different tilted angles. The traced subsystem $B$ with length $l=6$ is located at the left end of the system. The total system size is $L=64$ and the characteristic timescale $\tau_{1}$ is marked by vertical dashed line.
@@
where $\hat{n}_{j}=\hat{c}^{\dagger}_{j}\hat{c}_{j}$ is the on-site particle density. The results for $\overline{N}(t)$, exemplified in Fig. 1(a) for $\gamma=0.8$, reveal a universal long-time behavior. Here, “universal" refers to the robust relaxation towards half-filling ($L/2$) for any initial tilted ferromagnetic state that breaks the U(1) symmetry ($\theta>0$) within the Hatano-Nelson model under OBC. While we do not claim this extends to all non-Hermitian systems, we expect it to be a general feature for models exhibiting a prominent non-Hermitian skin effect.
This phenomenon is rigorously proven in the strong non-reciprocity limit. As we demonstrate analytically in section III-A of the SM [201] , for $\gamma=1$, the total particle number is guaranteed to asymptotically approach $L/2$. For weaker non-reciprocity ($\gamma<1$), while a full analytical proof is more challenging, our numerical results provide strong evidence for this trend. The clear convergence for $\gamma=0.8$ in Fig. 1(a) supports this. Furthermore, simulations presented in the SM [201] confirm that relaxation to half-filling still occurs for smaller $\gamma$, even under different boundary conditions for state preparation and evolution, showcasing the robustness of this phenomenon. This suggests that $\gamma$ primarily controls the rate of convergence to this universal steady state. In stark contrast, the number-conserving ferromagnetic state ($\theta=0$), being an eigenstate of the local density operator, remains fully occupied throughout the evolution. We also note that for the case of PBC, which is discussed in the SM [201] , the total particle number for a generic initial state also tends towards $L/2$ in the steady state.
@@
Density dynamics.— To further explore the dynamical characteristics reflected in the particle number, we now examine the time evolution of the on-site particle density $\langle\hat{n}_{j}\rangle$, exemplified in Fig. 1(b) for $\theta=\pi/6$ [201]. While the initial state exhibits a uniform particle density, distinct behaviors emerge at different timescales. At long times, there is a clear spatial separation of particles due to the skin effect. The particle density becomes concentrated on the left side of the system, leaving the opposite side sparsely populated. Intriguingly, short-time dynamics reveal a contrasting exotic behavior. Instead of accumulation, we observe an initial reduction of particle density near the left edge. This sharp depletion is linked to a chiral wavefront propagating with twice the maximum group velocity of quasiparticles. Specifically, particle density begins to decrease as the wavefront reaches a given site. Here, the group velocity is defined as $v_{\text{g}}=\frac{\mathrm{d}}{\mathrm{d}k}\text{Re}[E(k)]$, where $E(k)=-2J\text{cos}k-2i\gamma\text{sin}k$ is the energy spectrum of the HN model under PBC [204, 146, 205].
@@
With increasing time, two characteristic timescales, $\tau_{1}$ and $\tau_{2}$, become evident. The first $\tau_{1}$ corresponds to the arrival time of the chiral wavefront at the right boundary, at which point the total particle number approaches $L/2$ for large $\gamma$. Around this time, the particle density begins to change from both boundaries inwards: increasing from the left edge and decreasing from the right edge towards the center. Upon reaching the second timescale, $\tau_{2}$, the density distribution stabilizes into a steady state, with only minor oscillations remaining in the central region. This steady state distribution clearly demonstrates the skin effect, with the left half of the system almost fully occupied and the right half nearly empty. In stark contrast, an initial tilted antiferromagnetic (AFM) state lacks this chiral wavefront; its dynamics immediately resemble the post-$\tau_{2}$ behavior of the tilted ferromagnetic state, likely because the AFM state starts at half-filling. Detailed results are presented in section V of the SM [201].
@@
Similar to the particle density, characteristic timescales $\tau_{1}$ and $\tau_{2}$ are also evident in the dynamics of both current density and particle inflow. In the short-time regime ($\tau<\tau_{1}$), a chiral wavefront emerges in both the current density and particle inflow profiles. The characteristic time $\tau_{1}$ corresponds to the moment when this chiral wavefront reaches the right boundary of the system. Notably, the wavefront in the particle inflow in Fig. 1(d) represents a significant particle sink, indicating a substantial loss of particle number at given space-time location. The wavefront of current in Fig. 1(c) appears as a sharp spatial gradient of current density. Although the negative current gradient implies particle accumulation according to the continuity equation Eq. (8), the observed net effect at the wavefront is particle depletion. Thus, the particle number reduction associated with the wavefront in Fig. 1(b) arises from the effect of particle inflow acting as a propagating sink, which dominates over the particle-accumulating effect of the current induced by non-reciprocity. In the intermediate stage ($\tau_{1}<\tau<\tau_{2}$), the rapid increase in particle number at the left edge and decrease at the right edge are also driven by the particle inflow, as visualized in Fig. 1(d). For late times ($\tau>\tau_{2}$), both the current and particle inflow diminish, exhibiting only small, time-decaying oscillations. Results for other values of $\gamma$ show qualitatively consistent behaviors, as included in section V of the SM [201]. Moreover, our analysis of the density dynamics under APBC, detailed in the the section VI of the SM [201], yields qualitatively distinct conclusions, further demonstrating that the anomalous phenomena observed under OBC are attributed to the skin effect.
@@
Here, the system is bipartitioned into subsystems $A$ and $B$, and $\rho_{A}=\text{Tr}_{B}[\rho]$ is the reduced density matrix of subsystem $A$. For the Hermitian case (Fig. 2(a)), the EE for different initial tilted states initially increases and subsequently saturates to a steady-state value. In contrast, the non-Hermitian dynamics, shown in Fig. 2(b) and (c), exhibit more complex phenomena. Similar to the particle density dynamics, the entanglement evolution displays distinct behaviors in different temporal regimes. In the stage $\tau<\tau_{1}$, EE increases with evolution time and subsequently saturates, qualitatively consistent with the conventional behavior of entanglement dynamics in Hermitian systems. This trend is qualitatively captured by the quasiparticle picture, as demonstrated in section IV of the SM [201]. An intriguing observation, illustrated in Fig. 2(b)-(c), is that non-Hermiticity can enhance both the growth rate and the maximum value of EE during time evolution for specific initial states and small $\gamma$.
@@
Given the skin effect, we anticipate that the EE depends on the bipartition subsystem’s location. To investigate this, we calculate the EE for different subsystem positions, with detailed results presented in section V of the SM [201]. When the small subsystem is located centrally, the qualitative behavior of EE dynamics remains similar to the cases where the subsystem is at the left or right boundary. However, the reduction in EE after $t>\tau_{1}$ is significantly less pronounced compared to the results in Fig. 2(b) and (c). This milder reduction can be naturally understood through the particle number distribution induced by the skin effect, as the particle density fluctuation is stronger at the middle region.
@@
Recent research has revealed a novel and universal QME in dynamic symmetry restoration, where systems with greater initial symmetry breaking recover symmetry faster [156, 186, 183, 195, 185, 187, 188, 189, 190, 191, 192, 193, 194, 182]. A key question is whether symmetry restoration and the QME can emerge in non-Hermitian systems. Fig. 3 presents the second-order Rényi EA for varying non-Hermiticity strengths ($\gamma$). Consistent with prior work [156], Fig. 3(a) shows a clear curve crossing for different tilted angles ($\theta$) in the Hermitian case, confirming the QME. Notably, a pronounced QME also exists in the non-Hermitian scenario, as evidenced by curve crossings in Fig. 3(b) and (c). Intriguingly, this QME appears enhanced compared to the Hermitian case, with curve crossings occurring earlier in time. This enhancement is accompanied by non-monotonic behavior in the early-time EA evolution for some initial states, where EA initially increases before decreasing towards symmetry restoration. Further analysis reveals that the strength of this enhancement is itself non-monotonic with respect to the non-reciprocity strength $\gamma$. Our results suggest the existence of a critical value of $\gamma$ where the QME is most pronounced, before the effect weakens again at stronger non-reciprocity. A detailed discussion of this phenomenon is provided in the section V-E of the SM [201]. At late times, the EA results show a restoration of U(1) symmetry, mirroring the behavior observed in Hermitian systems.